# SocialFusion: Addressing Social Degradation in Pre-trained Vision-Language Models

### v1

@@ Abstract
Figure 1: SocialFusion achieves positive transfer across all visual social interaction understanding tasks. Popular open-source VLMs suffer from social degradation, where their visual encoders have been degraded by the visual-linguistic pre-training, leading to negative transfer when jointly trained on multiple visual social interaction understanding tasks. SocialFusion addresses this limitation by using a frozen visual encoder and a minimal fusion architecture, achieving positive transfer on all tested task metrics while remaining competitive across the board.
@@ 1 Introduction
Social interactions with family members, friends, colleagues, and strangers are an important component of people’s everyday lives. Understanding social interactions in visual media is a crucial capability for a visual intelligence system (lee2024socialaisurveyunderstanding) and has broad applications in augmented reality (geroimenko2023augmented), human–robot interactions (xu2021human), and assistive technologies (e.g., to help people with autism) (iannone2023breaking; robila2020applications).
Social interaction understanding comprises various visual tasks, as shown in Fig. 2. Prior work has identified five important categories made up of purely visual tasks (lee2024socialaisurveyunderstanding): gesture analysis, which helps to understand conversation (kendon1994gestures; de2019language); gaze and attention analysis, which disambiguates people’s intent and action (fathi2012learning; huang2015using; lukander2017inferring; emery2000eyes); facial expression analysis, which provides direct insight into emotion, intent, and wishes (horstmann2003facial); social situation analysis, inferring or reasoning about the social relationships, social norms, etc., in a scene; and conversation dynamics understanding, which involves tasks to understand speaker–addressee interactions, dialogue state tracking, etc. (lee2024socialaisurveyunderstanding).
Existing efforts toward solving such social interaction understanding tasks develop specialist models trained to solve one task at a time and feature task-specific architectures lee2024socialaisurveyunderstanding. While such models achieve good performance on many tasks, they are practically inconvenient for real-world use. A crucial step in building toward a general, socially competent system is to unify and jointly train diverse social interaction understanding tasks, especially if it is possible to achieve positive transfer by leveraging the synergies between different tasks.
General pre-trained vision-language models (VLMs) represent a natural candidate for unified social understanding, as they are known for a large number of diverse and general competencies (li2024survey; 10769058) due to their large-scale pre-training on similarly large and diverse datasets (hurst2024gpt; deitke2024molmo; wang2024qwen2; chen2025expandingperformanceboundariesopensource). We test several popular open-source VLMs of similar capacity: Qwen2-VL 2B (wang2024qwen2), Sail-VL 1.5 2B (dong-etal-2025-scalable), and MolmoE (deitke2024molmo), which is a mixture-of-experts model with 1B active parameters and 7B parameters in total. However, we find across the board that they are not able to foster consistent positive transfer between social tasks under parameter-efficient fine-tuning. In fact, as shown in Fig. 1, the models exhibit stark negative transfer, revealing an internal struggle to unify social competence.
@@ 2 Visual Social Interaction Understanding Tasks
We consider five visual social interaction understanding tasks to establish broad coverage over the task taxonomy presented in (lee2024socialaisurveyunderstanding). Details of each task are presented below.
@@
Gesture analysis.  Understanding human gestures and changes in body language is essential for interpreting social interactions (clough2020role; kendon1994gestures; de2019language). Therefore, a successful social understanding model must have the ability to classify different gestures. For this task, we use the HaGRIDv2 dataset, which has 33 different classes of gestures in over one million images (nuzhdin2024hagridv2), to classify gestures such as the “peace” gesture depicted in Fig. 2.
@@
Gaze and attention.  Gaze position estimation can provide insights about an individual’s intent and action (fathi2012learning; lukander2017inferring) and enables better understanding of social interactions (emery2000eyes), as eye movement has been shown to reflect the human thought process (yarbus2013eye). An example is shown in Fig. 2, where the man in the white shirt is focusing on the frisbee. The GazeFollow dataset (recasens2015they), containing 122,143 images with annotated gaze points, is a widely used benchmark for this task. Images with more than one person usually contain the gaze for each person. For that reason, the dataset also provides a bounding box annotation corresponding to each person. The GazeFollow test set has ten ground truth point annotations per sample from ten different annotators. The model’s prediction is compared against these ten annotations to find the minimum and average distances.
@@
Facial expression recognition.  It has been established in psychology that there are several fundamental human facial expressions that are universal across cultures (ekman1971constants; matsumoto1992more). These expressions can reveal insights on emotional state, behavioral intentions, and wishes (horstmann2003facial). An example with the “contempt” facial expression is shown in Fig. 2. In this paper, we use AffectNet (mollahosseini2017affectnet) for model training and evaluation, which contains 450,000 images manually annotated with eight emotional categories.
@@
Conversation dynamics analysis.  This task involves understanding the flow of a conversation by modeling aspects like who is speaking to whom, the targets of participants’ attention, and the evolving relationships of participants (lee2024socialaisurveyunderstanding). In this paper, we consider the “Looking at Me” (LAM) task from the Ego4D dataset (grauman2022ego4d). It classifies three million samples of faces of participants of a conversation based on whether they are currently looking at the camera wearer, which gives insight into the conversation members’ attention, demonstrated in Fig. 2 with a straightforward “Yes” classification.
@@
Social situation analysis.  This task includes understanding and inferring social relationships, answering social-reasoning questions, and other tasks that require an understanding of the social dynamic of a scene (lee2024socialaisurveyunderstanding). In the People in Social Context (PISC) benchmark (li2017dual), relationships are classified into predefined categories like “Friends,” as illustrated in Fig. 2. PISC has two subtasks: classifying the relationship among specific six classes (family, couple, friends, commercial, professional, and no relationship) and classifying the domain of the relationship among three broad classes (intimate, non-intimate, and no relationship). The dataset contains 76,568 samples over 22,670 images, where an image may contain more than two people and thus multiple relationship pair samples. For that reason, the dataset input contains another crucial element, a pair of bounding boxes for each relationship denoting the two individuals of interest.
@@
Together, these five tasks provide comprehensive coverage of visual social interaction understanding and enable unified modeling across diverse annotation types including classification labels, coordinate predictions, and bounding box inputs. The diversity of these tasks and their annotation formats presents both an opportunity for leveraging cross-task synergies and a challenge for unified modeling, which we address in the following section.
@@ 3.2 The Social Degradation Issue in Pre-Trained VLMs
We first experiment with three widely used VLMs of similar capacity: the 2B parameter variation of Qwen2-VL (wang2024qwen2), the 2B variation of Sail-VL 1.5 (dong-etal-2025-scalable), and for additional architectural variety, MolmoE (deitke2024molmo), a mixture-of-experts model with 1B active/7B total parameters.
We fine-tune each of these VLMs using a low-rank adapter (LoRA, hu2022lora) over the linear layers of the LLM module as a parameter-efficient and fair way to introduce new tasks. For each model, we consider two settings. We train each model both jointly, meaning training on all tasks simultaneously, and on each task separately. In joint training, we shuffle the datasets and randomly undersample them to the minimum size each epoch to avoid imbalanced training. Since we unified most tasks as plain-text generation, we are able to shuffle them within each batch. Otherwise, we randomly intersperse batches from the gaze heatmap prediction task with the regular text batches.
@@
We now begin to investigate why the pre-trained VLMs are not able to unify these social tasks in harmony. We hypothesize that the problem lies in the models’ visual encoders. In particular, we seek to measure of the impact of VLM pre-training on the encoders quantitatively. To this end, we compare each visual encoder before and after visual-linguistic pre-training. Qwen2-VL’s original visual encoder is DFN ViT (fang2023data), Sail-VL’s visual encoder is SailViT (dong-etal-2025-scalable), and MolmoE’s visual encoder is CLIP (radford2021learning).
@@
Table 2: VLM pre-training hurts visual encoders for gaze target estimation under the Gaze-LLE (ryan2025gazelle) architecture, under which we test the encoders of Qwen2-VL, MolmoE, and Sail-VL 1.5. All three encoders have degraded performance after VLM pre-training.
In this initial experiment, we begin with a single social task: gaze target estimation. We test the VLM encoders, with versus without VLM pre-training, in ryan2025gazelle’s Gaze-LLE architecture, which produces state-of-the-art accuracy using DINOv2 on the GazeFollow dataset.
@@ 3.3 Investigating VLM Social Degradation
Decodability.  We seek to understand the level of decodability of the embeddings produced by visual encoders before and after visual-linguistic pre-training. To that end, we use linear representation probes on the outputs of each encoder. Linear probing is a widely used method to investigate the inner workings of neural networks, allowing us to isolate the quality of each model’s learned representations for a specific task while minimizing architectural biases (saha2025pulseppgopensourcefieldtrainedppg; belinkov-2022-probing; alain2017understanding).
@@
$W^{*}_{t}=\operatorname*{arg\,min}_{W_{t}}\mathcal{L}(W_{t}(\operatorname{flatten}(f(X_{t})))+b_{t},Y_{t}),$
@@
Formally, let $\mathcal{L}_{t}(\theta)$ be the loss function for training the entirety of task $t\in\{1,\dots,T\}$ (as a single batch), parameterized by the weights $\theta$ of each encoder. Then, we can compute the gradient $g_{t}=\nabla_{\theta}\mathcal{L}_{t}(\theta)\in\mathbb{R}^{d}$ for all encoders on each task $t$. Following 10597963, we use the angle $\phi_{ij}$ between gradients $g_{i}$ and $g_{j}$, comparing training on task $i$ against task $j$, to compute the Gradient Conflict Degree (GCD) between each pair of encoders, defined as
@@ 3.4 Proposed Model
Unlike other VLMs, SocialFusion also supports 2D heatmap outputs indicating gaze positions by directly projecting and upscaling the final image features into a heatmap. Since some visual social tasks also include bounding boxes to refer to specific parts of the input, our model also natively accepts these as an additional input. Fig. 3 provides an overview of the model architecture. In the following, we introduce each component in detail.
Visual encoder.  To extract a feature-rich representation from visual input, SocialFusion relies on a large pre-trained foundation model. We primarily use CLIP (radford2021learning) as our visual encoder, which was trained to learn useful representations of images through natural language supervision. It has been successfully applied to diverse tasks in prior works (luddecke2022image; ryan2025gazelle; zhou2024image; zhang2022tip; zhao2023clip). We freeze the weights of the visual encoder to reduce the number of trainable parameters. To allow for fine-grained understanding of the visual input, we use the $336\times 336$ size input variant of CLIP and keep its weights frozen. With CLIP’s $14\times 14$ patch size, this results in a feature map $z\in\mathbb{R}^{24\times 24\times d_{v}}$ per image, where $d_{v}$ is the visual embedding dimension.
@@
Large language model (LLM).  The LLM of SocialFusion takes in a concatenated sequence [p,C(z)][\,p,C(z)\,], where $p$ is an embedded task-specific prompt that briefly describes the current task (see Appendix D for examples). This straightforward design promotes future work to similarly unify more multimodal inputs. For all tasks for which the ground truth can be represented in text, the model outputs the final model predictions as textual tokens, denoted by $\hat{y}$. In our experiments, we choose Llama 3.2 1B (grattafiori2024llama) as the LLM. As we did in the case of the visual encoder, we freeze all of the weights of the LLM. However, to allow it to learn to process and analyze the visual inputs, we add a small LoRA adapter (hu2022lora) to all of its linear layers, providing a fair comparison with the VLM baselines that we train in the same way.
Bounding box embedding.  In the gaze target estimation and social relationship recognition tasks, the input also contains bounding boxes denoting people of interest. Inspired by ryan2025gazelle, we incorporate bounding box information into the visual representations $z$ by adding a special learned position embedding, $p_{bbox}$, over features of patches that intersect with bounding boxes. Concretely, we construct a binary mask $M$ of the same dimension as the feature map $z$, where 1 indicates that a patch intersects with a bounding box and 0 means that it does not. Then, $z$ becomes
@@
Training Objective.  For the gaze heatmap output, the ground truth is a 2D Gaussian distribution (with $\sigma=3$) around the actual gaze point, following prior work (recasens2015they; chong2020detecting; ryan2025gazelle). We use a pixel-wise binary cross-entropy loss $\mathcal{L}_{heatmap}$ between the predicted and ground truth heatmaps. For all other tasks, since the ground truth is a sequence of textual tokens, we use a cross-entropy loss $\mathcal{L}_{llm}$ directly between the logits of the LLM and the ground truth tokens. The final loss is then

### v2

@@ Abstract
Figure 1: SocialFusion achieves positive transfer across all visual social interaction understanding tasks. Popular open-source VLMs suffer from social degradation, whereby their visual encoders have been degraded by the visual-linguistic pre-training, leading to negative transfer when jointly trained on multiple visual social interaction understanding tasks. SocialFusion addresses this limitation by using a frozen visual encoder and a minimal fusion architecture, achieving positive transfer on all tested task metrics while remaining competitive across the board.
@@ 1 Introduction
Social interactions with family members, friends, colleagues, and strangers are an important component of people’s everyday lives. Understanding social interactions in visual media is a crucial capability for a visual intelligence system (Lee et al., 2024b) and has broad applications in augmented reality (Geroimenko, 2023), human–robot interactions (Xu et al., 2021), and assistive technologies (e.g., to help people with autism) (Iannone and Giansanti, 2023; Robila and Robila, 2020).
Social interaction understanding comprises various visual tasks, as shown in Fig. 2. Prior work has identified five important categories made up of purely visual tasks (Lee et al., 2024b): gesture analysis, which helps to understand conversation (Kendon, 1994; De Stefani and De Marco, 2019); gaze and attention analysis, which disambiguates people’s intent and action (Fathi et al., 2012; Huang et al., 2015; Lukander et al., 2017; Emery, 2000); facial expression analysis, which provides direct insight into emotion, intent, and wishes (Horstmann, 2003); social situation analysis, inferring or reasoning about the social relationships, social norms, etc., in a scene; and conversation dynamics understanding, which involves tasks to understand speaker–addressee interactions, dialogue state tracking, etc. (Lee et al., 2024b).
Existing efforts toward solving such social interaction understanding tasks develop specialist models trained to solve one task at a time and feature task-specific architectures Lee et al. (2024b). While such models achieve good performance on many tasks, they are practically inconvenient for real-world use. A crucial step in building toward a general, socially competent system is to unify and jointly train diverse social interaction understanding tasks, especially if it is possible to achieve positive transfer by leveraging the synergies between different tasks.
General pre-trained vision-language models (VLMs) represent a natural candidate for unified social understanding, as they are known for a large number of diverse and general competencies (Li et al., 2024; Xu et al., 2025) due to their large-scale pre-training on similarly large and diverse datasets (Hurst et al., 2024; Deitke et al., 2024; Wang et al., 2024; Chen et al., 2025). We test several popular open-source VLMs of similar capacity: Qwen2-VL 2B (Wang et al., 2024), Sail-VL 1.5 2B (Dong et al., 2025), and MolmoE (Deitke et al., 2024), which is a mixture-of-experts model with 1B active parameters and 7B parameters in total. However, we find across the board that they are not able to foster consistent positive transfer between social tasks under parameter-efficient fine-tuning. In fact, as shown in Fig. 1, the models exhibit stark negative transfer, revealing an internal struggle to unify social competence.
@@ 2 Visual Social Interaction Understanding Tasks
We consider five visual social interaction understanding tasks to establish broad coverage over the task taxonomy presented in (Lee et al., 2024b). Details of each task are presented below.
@@
Gesture analysis.  Understanding human gestures and changes in body language is essential for interpreting social interactions (Clough and Duff, 2020; Kendon, 1994; De Stefani and De Marco, 2019). Therefore, a successful social understanding model must have the ability to classify different gestures. For this task, we use the HaGRIDv2 dataset, which has 33 different classes of gestures in over one million images (Nuzhdin et al., 2024), to classify gestures such as the “peace” gesture depicted in Fig. 2.
@@
Gaze and attention.  Gaze position estimation can provide insights about an individual’s intent and action (Fathi et al., 2012; Lukander et al., 2017) and enables better understanding of social interactions (Emery, 2000), as eye movement has been shown to reflect the human thought process (Yarbus, 2013). An example is shown in Fig. 2, where the man in the white shirt is focusing on the frisbee. The GazeFollow dataset (Recasens et al., 2015), containing 122,143 images with annotated gaze points, is a widely used benchmark for this task. Images with more than one person usually contain the gaze for each person. For that reason, the dataset also provides a bounding box annotation corresponding to each person. The GazeFollow test set has ten ground truth point annotations per sample from ten different annotators. The model’s prediction is compared against these ten annotations to find the minimum and average distances.
@@
Facial expression recognition.  It has been established in psychology that there are several fundamental human facial expressions that are universal across cultures (Ekman and Friesen, 1971; Matsumoto, 1992). These expressions can reveal insights on emotional state, behavioral intentions, and wishes (Horstmann, 2003). An example with the “contempt” facial expression is shown in Fig. 2. In this paper, we use AffectNet (Mollahosseini et al., 2017) for model training and evaluation, which contains 450,000 images manually annotated with eight emotional categories.
@@
Conversation dynamics analysis.  This task involves understanding the flow of a conversation by modeling aspects like who is speaking to whom, the targets of participants’ attention, and the evolving relationships of participants (Lee et al., 2024b). In this paper, we consider the “Looking at Me” (LAM) task from the Ego4D dataset (Grauman et al., 2022). It classifies three million samples of faces of participants of a conversation based on whether they are currently looking at the camera wearer, which gives insight into the conversation members’ attention, demonstrated in Fig. 2 with a straightforward “Yes” classification.
@@
Social situation analysis.  This task includes understanding and inferring social relationships, answering social-reasoning questions, and other tasks that require an understanding of the social dynamic of a scene (Lee et al., 2024b). In the People in Social Context (PISC) benchmark (Li et al., 2017), relationships are classified into predefined categories like “Friends,” as illustrated in Fig. 2. PISC has two subtasks: classifying the relationship among specific six classes (family, couple, friends, commercial, professional, and no relationship) and classifying the domain of the relationship among three broad classes (intimate, non-intimate, and no relationship). The dataset contains 76,568 samples over 22,670 images, where an image may contain more than two people and thus multiple relationship pair samples. The input therefore also contains a pair of bounding boxes denoting the relationship of interest.
@@
Together, these five tasks provide comprehensive coverage of visual social interaction understanding and enable unified modeling across diverse annotation types including classification labels, coordinate predictions, and bounding box inputs. The diversity of the tasks and their annotation formats presents both an opportunity for leveraging cross-task synergies and a challenge for unified modeling, which we address in the following section.
@@ 3.2 The Social Degradation Issue in Pre-Trained VLMs
We first experiment with three widely used VLMs of similar capacity: the 2B parameter variation of Qwen2-VL (Wang et al., 2024), the 2B variation of Sail-VL 1.5 (Dong et al., 2025), and for additional architectural variety, MolmoE (Deitke et al., 2024), a mixture-of-experts model with 1B active/7B total parameters.
We fine-tune each of these VLMs using a low-rank adapter (LoRA, Hu et al., 2022) over the linear layers of the LLM module as a parameter-efficient and fair way to introduce new tasks. For each model, we consider two settings. We train each model both jointly, meaning training on all tasks simultaneously, and on each task separately. In joint training, we shuffle the datasets and randomly undersample them to the minimum size each epoch to avoid imbalanced training. Since we unified most tasks as plain-text generation, we are able to shuffle them within each batch. Otherwise, we randomly intersperse batches from the gaze heatmap prediction task with the regular text batches.
@@
We now investigate why the pre-trained VLMs are not able to unify these social tasks in harmony. We hypothesize that the problem lies in the models’ visual encoders. In particular, we seek to measure of the impact of VLM pre-training on the encoders quantitatively. To this end, we compare each visual encoder before and after visual-linguistic pre-training. For Qwen2-VL, the original visual encoder is DFN ViT (Fang et al., 2023), for Sail-VL, it is SailViT (Dong et al., 2025), and for MolmoE, it is CLIP (Radford et al., 2021).
@@
Table 2: VLM pre-training hurts visual encoders for gaze target estimation under the Gaze-LLE (Ryan et al., 2025) architecture, under which we test the encoders of Qwen2-VL, MolmoE, and Sail-VL 1.5. All three encoders have degraded performance after VLM pre-training.
In this initial experiment, we begin with a single social task: gaze target estimation. We test the VLM encoders, with versus without VLM pre-training, in Ryan et al. (2025)’s Gaze-LLE architecture, which produces state-of-the-art accuracy using DINOv2 on the GazeFollow dataset.
@@ 3.3 Investigating VLM Social Degradation
Decodability.  We seek to understand the level of decodability of the embeddings produced by visual encoders before and after visual-linguistic pre-training. To that end, we use linear representation probes on the outputs of each encoder. Linear probing is a widely used method to investigate the inner workings of neural networks, allowing us to isolate the quality of each model’s learned representations for a specific task while minimizing architectural biases (Saha et al., 2025; Belinkov, 2022; Alain and Bengio, 2017).
@@
$W^{*}_{t}=\operatorname*{arg\,min}_{W_{t}}\mathcal{L}(W_{t}(\operatorname{flatten}(f(X_{t})))+b_{t},Y_{t}),\vskip-2.0pt$
@@
Formally, let $\mathcal{L}_{t}(\theta)$ be the loss function for training the entirety of task $t\in\{1,\dots,T\}$ (as a single batch), parameterized by the weights $\theta$ of each encoder. Then, we can compute the gradient $g_{t}=\nabla_{\theta}\mathcal{L}_{t}(\theta)\in\mathbb{R}^{d}$ for all encoders on each task $t$. Following Chai et al. (2024), we use the angle $\phi_{ij}$ between gradients $g_{i}$ and $g_{j}$, comparing training on task $i$ against task $j$, to compute the Gradient Conflict Degree (GCD) between each pair of encoders, defined as
@@ 3.4 Proposed Model
Unlike other VLMs, SocialFusion supports 2D heatmap outputs indicating gaze positions by directly projecting and upscaling the final image features into a heatmap. Since some visual social tasks also include bounding boxes to refer to specific parts of the input, our model also natively accepts these as an additional input. Fig. 3 provides an overview of the model architecture, which we introduce in detail in the following.
Visual encoder.  To extract a feature-rich representation from visual input, SocialFusion relies on a large pre-trained foundation model. We primarily use CLIP (Radford et al., 2021) as our visual encoder, which was trained to learn useful representations of images through natural language supervision. It has been successfully applied to diverse tasks in prior works (Lüddecke and Ecker, 2022; Ryan et al., 2025; Zhou et al., 2024; Zhang et al., 2022; Zhao et al., 2023). We freeze the weights of the visual encoder to reduce the number of trainable parameters. To allow for fine-grained understanding of the visual input, we use the $336\times 336$ size input variant of CLIP and keep its weights frozen. With CLIP’s $14\times 14$ patch size, this results in a feature map $z\in\mathbb{R}^{24\times 24\times d_{v}}$ per image, where $d_{v}$ is the visual embedding dimension.
@@
Large language model (LLM).  The LLM of SocialFusion takes in a concatenated sequence [p,C(z)][\,p,C(z)\,], where $p$ is an embedded task-specific prompt that briefly describes the current task (see Appendix D for examples). This straightforward design promotes future work to similarly unify more multimodal inputs. For all tasks for which the ground truth can be represented in text, the model outputs the final model predictions as textual tokens, denoted by $\hat{y}$. In our experiments, we choose Llama 3.2 1B (Grattafiori et al., 2024) as the LLM. As we did in the case of the visual encoder, we freeze all of the weights of the LLM. However, to allow it to learn to process and analyze the visual inputs, we add a small LoRA adapter (Hu et al., 2022) to all of its linear layers, providing a fair comparison with the VLM baselines that we train in the same way.
Bounding box embedding.  In the gaze target estimation and social relationship recognition tasks, the input also contains bounding boxes denoting people of interest. Inspired by Ryan et al. (2025), we incorporate bounding box information into the visual representations $z$ by adding a special learned position embedding, $p_{bbox}$, over features of patches that intersect with bounding boxes. Concretely, we construct a binary mask $M$ of the same dimension as the feature map $z$, where 1 indicates that a patch intersects with a bounding box and 0 means that it does not. Then, $z$ becomes
@@
Training Objective.  For the gaze heatmap output, the ground truth is a 2D Gaussian distribution (with $\sigma=3$) around the actual gaze point, following prior work (Recasens et al., 2015; Chong et al., 2020; Ryan et al., 2025). We use a pixel-wise binary cross-entropy loss $\mathcal{L}_{heatmap}$ between the predicted and ground truth heatmaps. For all other tasks, since the ground truth is a sequence of textual tokens, we use a cross-entropy loss $\mathcal{L}_{llm}$ directly between the logits of the LLM and the ground truth tokens. The final loss is then